# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ 4.4. Final step
Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, we know that Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$. By Step 2, the localization argument shows that Itô’s formula (2.3) holds for the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
@@
as $n\to\infty$. By Proposition 4.2, we have
@@
where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$. Now (4.17) follows from $\bm{X}\in{\mathcal{H}}^{p}$ and the dominated convergence theorem of stochastic integral ([37, Page 273, Lemma]). Next, since
@@
5. Proof of Section 3
5.1. Construction of the conditional independent copy
This subsection is devoted to the proof of the precise construction of conditional independent copy as in Theorem 3.1. For ease of exposition, we will show how to enlarge the probability space with two independent copies, and any finite number of copies can be easily adapted.
Recall the definition in Section 3, let us consider the probability space $\overline{\Omega}$ and the $\sigma$-algrebra $\overline{\mathcal{F}}$ defined by
$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ (5.1)
and probability measure $\overline{\mathbb{P}}$ is defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]},$ (5.2)
Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ is properly defined. Next, use the notation ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the nature extension into this enlarged probability space as
~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
@@ 5.1. Construction of the conditional independent copy
:=${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
With these constructions, we have the following propositions.
@@
Given $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ defined by (5.1) and (5.2) and the sub-algebras ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, ${\mathcal{F}}^{\prime}$, and ${\mathcal{F}}^{\prime\prime}$ defined by (5.3) and (5.4). Suppose $A,B\in{\mathcal{F}}$, let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$ and $C^{\prime\prime}=\Omega\times\Omega\times C$. Then
@@
    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$
    ${\mathbb{P}}(A|\mathcal{G})(\omega_{0})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})(\omega_{0},\omega_{1},\omega_{2}),\text{ a.s. under }\overline{\mathbb{P}}.$
@@
    Given $\widetilde{\mathcal{G}}$, then $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{\mathbb{P}}.$
    Moreover, this conditional probability ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ coincides with the conditional probability of $A^{\prime}$ and $A^{\prime\prime}$ given $\widetilde{\mathcal{F}}$, that is
    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}}),\text{ a.s. under }\overline{\mathbb{P}}.$
@@
    Given $\widetilde{\mathcal{G}}$, then $\widetilde{\mathcal{F}},{\mathcal{F}}^{\prime},$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent. That is,
    ¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s. under }\overline{\mathbb{P}}.$
    Moreover, given $\widetilde{\mathcal{F}}$, then ${\mathcal{F}}^{\prime}$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent, That is
    ¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{\mathbb{P}}.$
@@
To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is ${\mathcal{G}}$ measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}},\overline{\mathbb{P}})$ is $\widetilde{\mathcal{G}}$ measurable . For any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Recall the definition of $\overline{{\mathbb{P}}}$ at (5.2) and the property of conditional expectation, we have
¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
Again by (5.2), we obtain ¯$\overline{\mathbb{P}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{B}(\omega_{0})\overline{\mathbb{P}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
@@
¯$\overline{\mathbb{P}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]},$
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1). This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
As for the second statement of (2), note that ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})$ is $\widetilde{\mathcal{G}}$ measurable, thus $\widetilde{\mathcal{F}}$ measurable. For any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$. Then,
@@
¯$\overline{\mathbb{P}}(A^{\prime}\cap\widetilde{F})=\overline{\mathbb{P}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]},$
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1). This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ =¯$=\overline{{\mathbb{P}}}((A\cap G)\times B\times C)={\mathbb{E}}\Big{[}\mathbbm{1}_{A\cap G}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
Using the same calculation as in (5.5) and note that ${\mathbb{P}}(A|{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$, ${\mathbb{P}}(B|{\mathcal{G}})=\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})$ and ${\mathbb{P}}(C|{\mathcal{G}})=\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})$, we have
¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ $=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})\Big{]}.$
This leads to
¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{\mathbb{P}}.$
@@
As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$. Then,
¯$\overline{\mathbb{P}}(B^{\prime}\cap C^{\prime\prime}\cap\widetilde{F})$ =¯$=\overline{\mathbb{P}}(F\times B\times C)={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}$
@@
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{\mathbb{P}},$
which completes the whole proof. ∎
Proof of Theorem 3.1. For any random variable $R$ in $(\Omega,{\mathcal{F}},{\mathbb{P}})$, recall that the natural extension $\widetilde{R}$ and the two conditional independent copies $R^{\prime},R^{\prime\prime}$ of $R$ are defined by
@@
$\text{Law}\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=\text{Law}\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime\prime}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})\text{ a.s. under }\overline{\mathbb{P}},$
@@
$R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }\widetilde{\mathcal{G}}\text{ and }R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }{\mathcal{F}},$
which completes the proof of Theorem 3.1.∎
Proof of Corollary 3.2. It suffices to show that extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition, that is
~${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exists $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$ and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$. Then,
¯$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ =¯$=\overline{\mathbb{P}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$

### v3

@@ 4.4. Final step
Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$. By Step 2, the localization argument shows that Itô’s formula (2.3) holds for the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
@@
as $n\to\infty$. In fact, by Proposition 4.2, we have
@@
where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$. Now (4.17) follows from $\|\bm{X}\|_{{\mathcal{H}}^{p}}<\infty$ and Lemma 2.5.
Next, since
@@
5. Proof of Results in Section 3
5.1. Proof of Theorem 3.1
To prove the theorem, we first provide a constructive proof for the existence of conditionally independent copy claimed in Theorem 3.1. For ease of exposition, we will show how to enlarge the probability space with two independent copies, and the case with a finite number of copies can be easily adapted.
Recall the discussions in Section 3, and consider the probability space $\overline{\Omega}$, the $\sigma$-algebra $\overline{{\mathcal{F}}}$ defined by
$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\},$ (5.1)
and a probability measure $\overline{{\mathbb{P}}}$ defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]}.$ (5.2)
Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ is well defined.
Next, take $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the natural extension in this enlarged probability space as
$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
@@ 5. Proof of Results in Section 3
${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
In this framework, we have the following proposition.
@@
Consider $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ defined by (5.1) and (5.2) and the sub-algebras $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, $\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ defined by (5.3) and (5.4). Suppose $A,B\in{\mathcal{F}}$, and let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$, and $C^{\prime\prime}=\Omega\times\Omega\times C$. Then
@@
    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$,
[…]
@@
    Given $\widetilde{\mathcal{G}}$, $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
[…]
@@
    Given $\widetilde{\mathcal{G}}$, $\widetilde{\mathcal{F}}$$\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
    $\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
    Moreover, given $\widetilde{\mathcal{F}}$, then $\mathcal{F}$$^{\prime}$ and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
    $\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
@@
(of Proposition 5.1). To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is $\mathcal{G}$} measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is $\widetilde{\mathcal{G}}-$measurable. For any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. By the definition of $\overline{{\mathbb{P}}}$ in ((5.2)) and the property of conditional expectation,
$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
Again by (5.2), $\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{G}(\omega_{0})\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
@@
$\overline{{\mathbb{P}}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}$,
[…]
@@
$\overline{{\mathbb{P}}}(A^{\prime}\cap\widetilde{F})=\overline{{\mathbb{P}}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}$,
[…]
@@
To prove (3), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$, and
[…]
$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}.$
[…]
Therefore,
$\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{{\mathbb{P}}}.$
As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists an $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$, and by (5.2),
[…]
@@
where the the third equality follows from the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
$\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
Proof.
(of Theorem 3.1). For any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, recall that the natural extension $\widetilde{R}$ and the two conditionally independent copies $R^{\prime}$ and $R^{\prime\prime}$ of $R$ are defined by
@@
[…]
Remark 5.2.
Our construction is inspired by the proof of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]). Note that in these existing approaches, the common information $\mathcal{G}$} is modeled as part of the stochastic input. Moreover, in the context of weak and strong solutions of SDEs, one often uses the existence of regular conditional distributions to construct an extended probability space by embedding different weak solutions into a common space. To illustrate this, take the example of $(X^{(1)},W^{(1)})$ and $(X^{(2)},W^{(2)})$, let $\nu({\mathrm{d}}w)$ denote the common law of a Brownian motion, and define the regular conditional law $\eta^{(i)}({\mathrm{d}}x;w):=\text{Law}(X^{(i)}|W^{(i)}=w)$ for $i=1,2$. In such a setting, an extended probability measure is often defined (informally) by ${\mathbb{P}}({\mathrm{d}}x^{(1)}{\mathrm{d}}x^{(2)}{\mathrm{d}}w):=\eta^{(1)}({\mathrm{d}}x^{(1)};w)\cdot\eta^{(2)}({\mathrm{d}}x^{(2)};w)\nu({\mathrm{d}}w).$
Motivated by these constructions, we consider a more general setting in which the common information $\mathcal{G}$} may be abstract and regular conditional distributions may not be available. We define conditionally independent copies of some random variable $X$ directly by specifying the measure on the extended space supporting $X,X^{\prime}$ such that
${\mathbb{P}}\big{(}(X,X^{\prime})\in A\times B\big{)}:={\mathbb{E}}[\mathbbm{1}_{X\in A}\cdot{\mathbb{P}}(X\in B|{\mathcal{G}})],$
which remains consistent with the spirit of the classical constructions while avoiding the explicit use of regular conditional distributions.
Proof of Corollary 3.2. It suffices to show that the extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies
${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exist $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$, and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$. Then,
$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ $=\overline{{\mathbb{P}}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$